The small piston of a hydraulic lift has a diamete
Answer #1

pressure is distributed equally

force on small piston / area of small piston = force on big piston / area of big piston

force on small piston / pi * 4^2 = 15000 / pi * 20^2

force on small piston = 600 N

pressure = force / area

pressure = 600 / (pi * 0.04^2)

pressure = 119366.20 Pa